# Invariant Image Reparameterisation: A Unified Approach to Structural and Practical Identifiability and Model Reduction

### v1

@@ 1 Introduction
Establishing whether model parameters are uniquely determined is often referred to as structural identifiability analysis in the modelling literature [5, 16, 45, 50, 51, 25], and is also simply referred to as identifiability analysis in the statistical literature [46]. In contrast, learning whether sufficiently precise parameter estimates are possible when working with finite noisy observations corresponds to the notion of practical identifiability (or estimability) analysis [33, 46, 70, 41].
@@
Parameter non-identifiability occurs when different combinations of parameters produce identical, or sufficiently close, model outputs. making it impossible to uniquely determine the underlying mechanisms from observations [46, 65].
When model parameters are not identifiable, key questions arise: Can we identify certain parameter combinations even when individual parameters are non-identifiable [46, 17]? Can we perform model reduction to obtain simpler, identifiable models [13, 14, 27, 50, 51, 17]? Can we use practical non-identifiability to guide experimental design [63] or motivate approximate model reduction [49]?
@@
Existing approaches include symbolic methods [13, 14, 18, 15], profile likelihood analysis [40, 60, 49, 64, 68], and sloppiness analysis [8, 31, 53]. Related work includes active subspaces [19] and likelihood-informed dimension reduction [23, 24] for large-scale inverse problems. While each approach has strengths, they also have limitations: symbolic methods typically apply only to idealised data/structural identifiability scenarios, profile likelihood methods usually require manual choice of target parameter to analyse, and sloppiness analysis may use non-parameterisation-invariant approximations for uncertainty quantification. Furthermore, the relationship between sloppiness and identifiability is not always clear, as sloppiness is a relative notion quantified in terms of ratios of eigenvalues (or singular values) models can be sloppy yet identifiable [16]. In applications of sloppiness and profile likelihood analysis to model reduction, users often simply set poorly-identified parameters to zero or some arbitrary value [42, 26, 69]. However, this can lead to setting individually poorly-identified parameters to values that jointly violate the requirements on possible values for the well-identified parameter combinations. Model reduction methods for inverse problems [19, 23], while powerful for computation, particularly Bayesian inverse problens, do not typically directly address identifiability questions and do not usually aim to provide interpretable parameter combinations for mechanistic models.
@@
We present a novel approach that: (i) automatically identifies well-identified parameter combinations without symbolic computation; (ii) enables model reduction without requiring an explicit reduced form model; and (iii) provides statistically sound interval estimates for both parameters and predictions. Our method combines ideas from symbolic analysis and sloppiness literature to convert the general symbolic problem to a numerical one solvable at a single reference point. These parameter combinations then serve as interest parameters in our Profile-Wise Analysis (PWA) framework [65, 64]. For certain parameter combination classes, we recover structural identifiability analysis results to arbitrary accuracy, while still discovering useful approximate combinations in other cases.
@@ 2.1 Terminology
We work relatively informally with (assumed-to-be) smooth mappings between finite-dimensional spaces [44], using standard terminology for submersions – mappings with full rank derivative – and immersions – mappings with injective derivative. The observed Fisher information measures parameter information content via the negative Hessian of the log-likelihood [58, 20, 59]. Informally, parameters or parameter combinations are identifiable if they can be uniquely determined from ‘perfect data’. In the context of a model mapping parameters to data distributions, $\theta\mapsto p(y;\theta)$, this is equivalent to the mapping, or more generally an induced relation in terms of parameter combinations, being injective [46].
@@ 2.2 Model representation
For spatial/temporal models, we use numerical solutions on a fine grid as the ‘exhaustive’ summary, evaluating derivatives via automatic differentiation in Julia [61]. Thus, while our approach is essentially identical to the ‘exhaustive summary’ approach reviewed by [18, 17], we generally avoid the need for explicit exhaustive summary methods such as Taylor series or generating series (as reviewed in the cited articles), though these can be used if available. In this context, we emphasise that automatic differentiation at the source code level is distinct from symbolic differentiation and only evaluates the numerical value of the derivative at a fixed input value [30].
@@ 2.4 Symbolic reparameterisation condition
Given null space vectors $\alpha(\theta)$ obtained from $D_{\theta}\phi(\theta)$, we can then solve:
@@
for $\psi(\theta)$. This is equivalent to the conditions given by e.g. [12, 13], summarised by  [18, 17]. It’s also equivalent to equating the rows of $D_{\theta}\psi(\theta)$ to the vectors spanning the row space of $D_{\theta}\phi(\theta)$. The image of these vectors under $\psi$ gives the intermediate ‘image’ component $C$ of the epi-mono factorisation. As shown in detail in the supplementary material, we can also work directly with the observed Fisher information by solving a null space constraint in terms of the Fisher information rather than the Jacobian derivative of the auxiliary mapping. When we use the limiting Fisher information (i.e., as evaluated at the discretisation level of the solution grid), the approaches are equivalent; if not, the Fisher information formulation also includes additional finite data effects (see below).
@@ 2.5 Parameter space linearisation
Working in log coordinates, defining $\theta^{*}=\log(\theta)$ to be the component-wise log of the parameter vector, models that can be written in terms of monomial parameter combinations take the form:
(10) $\phi(\theta)=\tilde{\phi}(\exp(A\theta^{*})),$
where $A$ is a matrix of real coefficients. The chain rule gives:
@@
This follows since the first term on the right-hand side of the chain rule equation (11) is injective by assumption, and the second term is a diagonal matrix of (non-zero) exponentials and hence also injective. As before (see also supplementary material), these leave the null space of $A$ unchanged. As the left-hand side of (12) is parameter-independent, the right-hand side must be too. We can hence evaluate the right-hand side derivative at any convenient point to determine the global, invariant null space, complementary row space and hence intermediate image, reducing what would be a PDE to an algebraic computation. If the model is assumed non-identifiable, the point used will generally be, e.g., any (non-unique) maximum likelihood estimate, and the Jacobian will be singular. This is expected and causes no issues in general.
While this condition does not uniquely determine $A$, as any basis for the null space (and complementary row space) can be used, it establishes the existence of a matrix capturing non-identifiable and identifiable parameter combinations. While the coefficients in $A$ are traditionally restricted to integers in dimensional analysis, here we allow them to be arbitrary real numbers and do not require the combinations to be dimensionless a priori. This relaxation facilitates numerical computation and provides more flexibility in discovering practically useful parameter combinations. However, we also consider a posteriori scaling and rounding operations to convert these real-valued coefficients to integers to aid interpretability while preserving key properties of the approximation.
@@ 2.6 Singular value decomposition for nonlinear parameter combinations
(13) $D_{\theta^{*}}\phi_{*}(\hat{\theta}^{*})=U\Sigma V^{T}=U_{r}\Sigma_{r}V_{r}^{T}$
Taking $A=V_{r}^{T}$, or any row-scaled version of this, ensures the same kernel as $D_{\theta^{*}}\phi_{*}(\hat{\theta}^{*})$. The singular values provide a natural ranking of parameter combinations by identifiability, with zero or near-zero values indicating structural or practical non-identifiability respectively.
We can also partition parameters into identifiable and non-identifiable combinations using the full rather than reduced SVD:
(14) $\theta\mapsto\begin{bmatrix}\psi(\theta)\\ \lambda(\theta)\end{bmatrix}=\exp(V^{T}\log(\theta))$
where $\psi(\theta)$ corresponds to non-zero singular values and associated singular vectors, and $\lambda(\theta)$ to (near) zero values and associated singular vectors. This maintains compatibility with existing model codes while capturing the identifiability structure.
@@
As mentioned above, the null space and row space of the limiting observed Fisher information are the same as that of the auxiliary mapping, we can also extend this approach to work directly with the observed Fisher information (supplementary information). This may be more natural for many models.
@@ 2.7 Practical vs structural identifiability and observation operators
An important aspect of the relationship between practical and structural identifiability is represented by the need for observation operators such as:
@@
where $B_{\text{obs}}$ maps the fine-scale (numerical) solution to (experimental) observation points. The additional loss of information due to coarser observation scale can introduce additional null space components or poorly-identified parameters depending on the degree of additional discretisation. When we discuss the limiting observed Fisher information, we mean the observed Fisher information at the solution grid level; this is required for the Fisher information analysis to match the structural identifiability analysis, otherwise additional null space/poorly-identified components may be introduced.
When analysing at the solution grid level, our results should, to numerical approximation, recover those from an idealised structural identifiability analysis. However, for ill-posed problems or models with parameter-dependent limiting behavior, the distinction between practical and structural non-identifiability can blur [46]. The auxiliary mapping may have arbitrarily small singular values in certain parameter regions, making the distinction between practical and structural non-identifiability less clear. Furthermore, it is unclear if ideal structural identifiability results are useful in the absence of practical identifiability [46]. Our approach here is to use the auxiliary mapping (or limiting Fisher information) to determine the reparameterisation, but carry out the likelihood-based uncertainty analysis at the observation grid level. This provides a practical identifiability analysis that is consistent with the structural identifiability analysis
@@ 2.9 Profile-wise analysis
We quantify uncertainty using our PWA framework [64], which connects identifiability, parameter estimation, and predictive uncertainty through likelihood-based confidence sets. This approach uses both the joint likelihood for all parameters and profiles of the likelihood – the likelihood as a function of a target parameter with the other ‘nuisance’ parameters maximised out [40, 60, 20, 59]. As the likelihood function is invariant to reparameterisation [58, 59, 20], reparameterisation of the full likelihood preserves all information in the original likelihood, unlike quadratic approximations based on the Fisher information. Rather, reparameterisation presents the information in the likelihood in a clearer form, while profile likelihoods reveal identifiability of individual parameters, or parameter combinations, separately from nuisance parameters [2].
@@ 4.3 Flow in heterogeneous media
(a) Joint likelihood in $(T_{1},T_{2})$ plane.
(b) Joint likelihood in $(T_{1},R)$ plane.
(c) Joint likelihood in $(T_{2},R)$ plane.
@@
(g) Joint likelihood in $(\frac{T_{2}}{R},\frac{T_{1}}{\sqrt{T_{2}R}})$ plane.
(h) Joint likelihood in $(\frac{T_{1}}{\sqrt{T_{2}R}},T_{1}T_{2}R)$ plane.
(i) Joint likelihood in $(\frac{T_{2}}{R},\frac{T_{1}}{\sqrt{T_{2}R}})$ plane.
@@
5 Conclusion and Future Work
Here we have presented Invariant Image Reparameterisation (IIR), an approach to parameter identifiability analysis that bridges structural and practical identifiability analysis, and symbolic and numerical approaches to implementation. We show how an initial coordinate transformation converts the symbolic approach of discovering identifiable parameter combinations into numerical computation at a single reference point. This numerical approach recovers symbolic results for a common class of parameter combinations while providing a natural ranking of such parameter combinations by their degree of identifiability in general and in cases of practical non-identifiability.
We find that practical non-identifiability, particularly when one-sided or asymmetric, is closely related to the existence of potential model reductions, mirroring observations by [49] on using profile likelihood for model reduction. Our use of profile-likelihood-based Profile-Wise Analysis for uncertainty quantification, in addition to Fisher information-based reparameterisation, reveals such asymmetric limits in terms of likelihood contours and profiles, in contrast to approaches focusing solely on Fisher-information/Hessian analysis.
As in [64], our approach extends to more complex models such as stochastic differential equations or simulation-based models, given an appropriate auxiliary mapping. Here, we used the full model solution as our ‘exhaustive summary’, while stochastic or simulation-based models generally need more carefully defined data distribution parameters such as moment dynamics equations [9]. Additionally, while our approach applies in principle to larger models, we deliberately considered examples with low parameter dimension, as even these give complex results. Further work could explore applications to larger models and spatially/temporally varying parameters, as well as improving scaling and rounding heuristics for interpretable parameter combinations.
The provided Julia implementation makes our methodology accessible for both analysis and further development. We believe IIR will prove valuable for model development and analysis across the quantitative sciences, particularly in complex applications where symbolic approaches may be impractical.

### v2

@@ 1 Introduction
Establishing whether model parameters are uniquely determined by ideal or ‘infinite’ data is often referred to as structural identifiability analysis in the modelling literature [5, 17, 47, 52, 53, 26], and is also simply referred to as identifiability analysis in the statistical literature [48]. In contrast, learning whether sufficiently precise parameter estimates are possible when working with finite noisy observations corresponds to the notion of practical identifiability (or estimability) analysis [35, 48, 72, 43].
@@
Parameter non-identifiability occurs when different combinations of parameters produce identical, or sufficiently close, model outputs, making it impossible to uniquely determine the underlying mechanisms from even perfect observations of the outputs [48, 67].
@@
When model parameters are not identifiable, key questions arise: Can we identify certain parameter combinations even when individual parameters are non-identifiable [48, 18]? Can we perform model reduction to obtain simpler, identifiable models [14, 15, 29, 52, 53, 18]? Can we use practical non-identifiability to guide experimental design [65] or motivate approximate model reduction [51]? Various approaches to addressing such questions include symbolic methods [14, 15, 19, 16], profile likelihood analysis [42, 62, 51, 66, 70], and sloppiness analysis [9, 33, 55]. Related work in large-scale inverse problems includes active subspaces [20, 7] and likelihood-informed dimension reduction [24, 25]. While each approach has strengths, they also have limitations: symbolic methods typically apply only to idealised data/structural identifiability scenarios, profile likelihood methods usually require user choice of target parameter to analyse, and sloppiness analysis may use non-parameterisation-invariant approximations for uncertainty quantification.
While systematic approaches to parameter subset selection for profile likelihood analysis have been developed, e.g. using Fisher Information Matrix analysis combined with profile likelihood [27] or context-dependent analysis along profiles [51], these approaches still require careful parameter subset selection steps and determination of functional relationships from profiles. The connection between sloppiness and identifiability is also somewhat unclear, as sloppiness is a relative notion quantified in terms of ratios of eigenvalues (or singular values) and models can be sloppy yet identifiable [17]. In applications of sloppiness and profile likelihood analysis to model reduction, users often simply set poorly-identified parameters to zero or some arbitrary value [44, 28, 71]. However, this can lead to setting individually poorly-identified parameters to values that jointly violate the requirements on possible values for the well-identified parameter combinations. Model reduction methods for inverse problems [20, 24], while powerful for computation, particularly Bayesian inverse problems, do not typically directly address identifiability questions (though see [7]) and do not usually aim to provide interpretable parameter combinations for mechanistic models.
@@
We present a novel approach that: (i) automatically identifies well-identified parameter combinations without symbolic computation; (ii) enables model reduction without requiring an explicit reduced form model; and (iii) provides statistically sound interval estimates for both parameters and predictions. Our method combines ideas from symbolic analysis and sloppiness literature to convert the general symbolic problem to a numerical one solvable at a single reference point. These parameter combinations then serve as interest parameters in our Profile-Wise Analysis (PWA) framework [67, 66]. For models expressible as functions of monomial parameter combinations, we recover structural identifiability analysis results to arbitrary accuracy, while still discovering useful approximate monomial combinations in other cases. Our approach also allows for non-integer powers in the ‘monomial’ combinations.
To illustrate these ideas, we first apply our method to a simple statistical model with practical identifiability issues and an approximate limiting model, before considering mechanistic models based on differential equations.
@@ 2.1 Terminology
We work relatively informally with (assumed-to-be) smooth mappings between finite-dimensional spaces [46], using standard terminology for submersions – mappings with full rank derivative – and immersions – mappings with injective derivative. The observed Fisher information measures local parameter information content via the negative Hessian of the log-likelihood [60, 21, 61], with zero eigenvalues indicating parameter redundancy and potential non-identifiability [13]. While we mention the Fisher information as a complementary approach to detecting parameter redundancy, our primary focus is on working directly with the auxiliary mapping and likelihood. Informally, parameters or parameter combinations are identifiable if they can be uniquely determined from ‘perfect data’. In the context of a model mapping parameters to data distributions, $\theta\mapsto p(y;\theta)$, this is equivalent to the mapping, or more generally an induced relation in terms of parameter combinations, being injective [48].
@@ 2.2 Model representation
For spatial/temporal models, we use numerical solutions on a fine grid as the ‘exhaustive’ summary, evaluating derivatives via automatic differentiation in Julia [63]. Thus, while our approach is essentially identical to the ‘exhaustive summary’ approach reviewed by [19, 18], we generally avoid the need for explicit exhaustive summary methods such as Taylor series or generating series (as reviewed in the cited articles), though these can be used if available. In this context, we emphasise that automatic differentiation at the source code level is distinct from symbolic differentiation and only evaluates the numerical value of the derivative at a fixed input value [32]. Numerical solutions are typically more readily available and applicable than symbolic methods for complex models.
@@ 2.4 Symbolic reparameterisation condition
Given null space vectors $\alpha(\theta)$ obtained from $D_{\theta}\phi(\theta)$, we can then solve the differential condition:
@@
for $\psi(\theta)$. This is equivalent to the conditions given by e.g. [13, 14], summarised by  [19, 18]. It is also equivalent to equating the rows of $D_{\theta}\psi(\theta)$ to the vectors spanning the row space of $D_{\theta}\phi(\theta)$, which gives a system of differential equations for $\psi$. The image of the parameter space $\Theta$ under the resulting $\psi$ gives the intermediate ‘image’ component $C$ of the epi-mono factorisation. A concrete example of solving the above symbolic condition is provided in the Results section.
As shown in the supplementary material, we can also work directly with the observed Fisher information by solving a null space constraint in terms of the Fisher information rather than the Jacobian derivative of the auxiliary mapping. For spatial or temporal models, we need to account for the functional nature of the model outputs. While we adopt a discretisation approach to representing model outputs, here the discretisation level used – solution grid or observation grid – matters. In particular, the auxiliary mapping and the observed Fisher information must be defined on the same grid level for the null spaces to match, otherwise additional null space components may be introduced (see supplementary material and discussion below). As we work with the auxiliary mapping defined on the solution grid level for determining the reparameterisation, this is only relevant for potential comparisons to Fisher information-based methods.
@@ 2.5 Parameter space linearisation
To work in log coordinates, we define $\theta^{*}=\log(\theta)$ to be the componentwise log of the parameter vector, and define $\phi_{*}(\theta^{*})=\phi(\theta)$ and $\psi_{*}(\theta^{*})=\psi(\theta)$ to be the corresponding mappings in log coordinates. Models that can be written in terms of monomial parameter combinations take the form:
(10) $\phi(\theta)=\phi_{*}(\theta^{*})=\tilde{\phi}(\exp(A\theta^{*})),$
where $A$ is a matrix of coefficients and $\tilde{\phi}$ is the intermediate mapping in (5). The chain rule gives:
@@
This follows since the first term on the right-hand side of the chain rule equation (11) is injective by assumption, and the second term is a diagonal matrix of exponential entries (which are hence non-zero) and is therefore injective. As before (see also supplementary material), these leave the null space of $A$ unchanged. As the left-hand side of (12) is parameter-independent, the right-hand side must be too. We can hence evaluate the right-hand side derivative at any convenient point to determine the global, invariant null space, complementary row space and hence intermediate image, reducing what would be a PDE to an algebraic computation. If the model is assumed non-identifiable, the point used will generally be, e.g., any (non-unique) maximum likelihood estimate, and the Jacobian will be singular. This is expected and causes no issues in general.
While the above condition does not uniquely determine $A$, as any basis for the null space (and complementary row space) can be used, it establishes the existence of a matrix capturing non-identifiable and identifiable parameter combinations. While the coefficients in $A$ are traditionally restricted to integers in dimensional analysis, here we allow them to be arbitrary real numbers and do not require the combinations to be dimensionless a priori. This relaxation facilitates numerical computation and provides more flexibility in discovering practically useful parameter combinations. However, we also consider a posteriori scaling and rounding operations to convert these real-valued coefficients to integers to aid interpretability while preserving key properties of the approximation.
@@ 2.6 Singular value decomposition for nonlinear parameter combinations
(13) $D_{\theta^{*}}\phi_{*}(\hat{\theta}^{*})=U\Sigma V^{T}=U_{r}\Sigma_{r}V_{r}^{T}.$
Taking $A=V_{r}^{T}$, or any row-scaled version of this, ensures the same kernel as $D_{\theta^{*}}\phi_{*}(\hat{\theta}^{*})$. The singular values provide a natural ranking of parameter combinations by identifiability. Zero or near-zero values indicate structural or practical non-identifiability, respectively, corresponding to insensitivity of the auxiliary mapping to these parameter combinations. The associated singular vectors provide the associated parameter combinations. The SVD is chosen because it is generally numerically stable and provides an orthogonal basis for the parameter space, maximising the separation of information concerning the different parameter combinations (see also [2, 22]).
We can also partition parameters into identifiable and non-identifiable combinations using the full, rather than reduced, SVD:
(14) $\theta\mapsto\begin{bmatrix}\psi(\theta)\\ \lambda(\theta)\end{bmatrix}=\exp(V^{T}\log(\theta)),$
where $\psi(\theta)$ corresponds to non-zero singular values and associated singular vectors, and $\lambda(\theta)$ to (near) zero values and associated singular vectors. This maintains compatibility with existing model codes that require full parameter vectors, while capturing the identifiability structure.
@@ 2.7 Practical vs structural identifiability and observation operators
An important aspect of the relationship between practical and structural identifiability for spatial or temporal models is represented by the need for observation operators, such as:
@@
where $B_{\text{obs}}$ maps the fine-scale (numerical) solution to (experimental) observation points. The additional loss of information due to coarser observation scale can introduce additional null space components or poorly-identified parameters depending on the degree of additional discretisation (see supplementary material).
When analysing at the solution grid level, our results should, to numerical approximation, recover those from an idealised structural identifiability analysis. However, for ill-posed problems or models with parameter-dependent limiting behavior, the distinction between practical and structural non-identifiability can blur [48]. The auxiliary mapping may have arbitrarily small singular values in certain parameter regions, making the distinction between practical and structural non-identifiability less clear. Furthermore, it is unclear if ideal structural identifiability results are useful in the absence of practical identifiability [48]. Our approach here is to use the auxiliary mapping at the solution grid level to determine the reparameterisation based on the full SVD in log coordinates. This reparameterisation allows us to group parameters according to degree of identifiability rather than strict identifiability or non-identifiability, making the distinction between practical and structural identifiability is less critical for our approach. However, given a parameterisation, we carry out the likelihood-based uncertainty analysis at the observation grid level to further reflect the practical identifiability of the model. Thus we provide a practical identifiability analysis that is consistent with, but not reliant on, idealised structural identifiability analysis.
@@ 2.9 Profile-wise analysis
We quantify uncertainty using our PWA framework [66], which connects identifiability, parameter estimation, and predictive uncertainty through likelihood-based confidence sets. This approach uses both the joint likelihood for all parameters and profile likelihoods, where the likelihood is evaluated as a function of a target parameter while maximising over the other ‘nuisance’ parameters for each value of the target parameter [42, 62, 21, 61]. We implement our methods numerically in Julia, with full details, including the packages used, available in the repository.
As the likelihood function is invariant to reparameterisation [60, 61, 21], reparameterisation of the full likelihood preserves all information in the original likelihood, unlike quadratic approximations based on the Fisher information. Rather, reparameterisation presents the information in the likelihood in a clearer form, while profile likelihoods reveal identifiability of individual parameters, or parameter combinations, separately from nuisance parameters [2].
@@ 4.3 Flow in heterogeneous media
(a) Joint profile likelihood in $(T_{1},T_{2})$ plane.
(b) Joint profile likelihood in $(T_{1},R)$ plane.
(c) Joint profile likelihood in $(T_{2},R)$ plane.
@@
(g) Joint profile likelihood in $(\frac{T_{2}}{R},\frac{T_{1}}{\sqrt{T_{2}R}})$ plane.
(h) Joint profile likelihood in $(\frac{T_{1}}{\sqrt{T_{2}R}},T_{1}T_{2}R)$ plane.
(i) Joint profile likelihood in $(\frac{T_{2}}{R},\frac{T_{1}}{\sqrt{T_{2}R}})$ plane.
@@
5 Conclusions and Future Work
Here we have presented Invariant Image Reparameterisation (IIR), an approach to parameter identifiability analysis that bridges structural and practical identifiability analysis and symbolic and numerical methods of implementation. We show how an initial coordinate transformation can convert the symbolic approach of discovering identifiable parameter combinations into numerical computation at a single reference point. This numerical approach recovers symbolic results for a typical class of parameter combinations while providing a natural ranking of such parameter combinations by their degree of identifiability in general and in cases of practical non-identifiability. While IIR reduces reliance on symbolic methods, these remain useful in cases where the parameter combinations of interest are more complex than monomials and when the model is amenable to symbolic computation. Extending the present work to such cases is of interest for future work. Relatedly, it would be of interest to investigate the possibility of extending the parameter transformation approach of IIR to rational functions of parameters, as used for fitting to profile relationships in [27]. We emphasise, however, that monomial parameter combinations are common in practice based on dimensional considerations and asymptotic arguments connected to model reduction [3].
Relatedly, we find that practical non-identifiability, particularly when one-sided or asymmetric, is closely related to the existence of potential model reductions when a parameter combination becomes small or large, mirroring observations by [51] on using profile likelihood for model reduction. Our use of profile-likelihood-based Profile-Wise Analysis for uncertainty quantification, in addition to reparameterisation based on the auxiliary mapping (or Fisher information), reveals such asymmetric limits in terms of likelihood contours and profiles, in contrast to approaches focusing solely on Fisher-information/Hessian analysis.
As in [66], our framework in principle extends to more complex models such as stochastic differential equations or simulation-based models, given an appropriate auxiliary mapping. However, while here we used the full model solution as our ‘exhaustive summary’, stochastic or simulation-based models generally need more carefully defined data distribution parameters such as moment dynamics equations [10]. Additionally, while our approach applies in principle to larger models, we deliberately considered examples with low parameter dimension, as even these give complex results. Further work could explore applications to larger models and spatially/temporally varying parameters, as well as improving scaling and rounding heuristics for interpretable parameter combinations.
The provided Julia implementation makes our methodology accessible for future work and development. We believe IIR will prove valuable for modelling and data analysis across the quantitative sciences, particularly in complex applications where symbolic approaches may be impractical.